[Hallc_running] [New Logentry] RC report for Sat. 20 and Sunday 21st of July

mboer at jlab.org mboer at jlab.org
Sun Jul 21 20:00:03 EDT 2019

Logentry Text:
RC report, Sat. 20 and Sunday 21st of July

Please find below RC notes. Sorry for not having posted reports earlier during the week end. Things are running more smoothly now.

- Saturday morning: 

End of the pion electroproduction experiment E12-06-101/E12-07-105 at 8AM. We switched to the VCS experiment E12-15-001. We had a controlled access to the hall at 8AM. Jack and Andy went down to switch HMS polarity from negative to positive. Brad lower down the hgc pressure of C4F8O to .35 atm. 

We were ready to take data around 11AM. We started with tuned beam, then with elastic runs single arm and coin to set trigger timing. 

New trigger setting is : 
PS1= HMS 3/4, PS4= SHMS 3/4, PS5= SHMS EL-real / HMS 3/4, PS6 = SHMS el-real / HMS 3/4. We intend to use PS6 for physics coincidence runs. Brad adjusted the timing. 

Data reconstruction seems fine, however we observe 2 coincidence peaks in elastic data which can be from trigger window. 

- Saturday afternoon:

We started the first VCS physics run at 5PM. coin=PS6. Observed high rates at 20 uA, from elastic. Mark disable hodo plans to get rid of part of elastics and Brad adjusted HMS 3/4  and SHMS 3/4  timing gates. 

We took good data and things went smoothly for the rest of the afternoon. However, rates stayed high and we kept the beam intensity between 15 uA to 25 uA. We still observe 2 peaks in coincidence time. Several good runs at VCS physics kinematics Ia and Ib were taken. 

- Saturday to Sunday night:

While doing rotation to move to kinematic Ia (HMS 51.5°) to IIa (HMS 33.1°) the drive controller striped. Steve Wood went down to the hall on restricted access to try to reset HMS and SHMS drivers, but it was not possible. We called PD to ask for a radcon survey to have Steve Lassiter work in the hall Sunday morning.
HMS got stuck at an angle close to the goal, we decided to stay at that angle (33.5°), take data for the night and update kinematics and run plan accordingly. We ran smoothly during the night.

- Sunday morning and afternoon:

Radcon performed a radiation survey and Steve went to work in the hall. We were good to take beam at 11:30AM. Started to run at the next kinematic setting, matching the one we ran during the night. Noticed at 7PM that momentum was properly updated, not angle (but small offset, data are usable, it will be taken into account). We rotated HMS accordingly and decided to keep running at this kinematics until 8AM to not take the risk of getting stuck again on Sunday night. Everything is running well so far.

Mark figured out the problem in coincidence time (at software level), he is working on it.

- Next:

We will change our kinematic settings (HMS) at the shift change tomorrow morning. 
We will have a RC meeting tomorrow at 4:30 PM.



This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3708037
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190721/7610e838/attachment.html>

More information about the Hallc_running mailing list